OpenStudy (anonymous):

let v1=(3,5) and v2=(-4,7). Compute |v1| and |v2|.

jimthompson5910 (jim_thompson5910):

if v = <x,y> then |v| = sqrt( x^2 + y^2 )

OpenStudy (anonymous):

So then we would plug in the numbers for each one to get it to equal the |v1| and |v2|

jimthompson5910 (jim_thompson5910):

In the case of v1 = <3, 5>, we get |v| = sqrt( x^2 + y^2 ) |v1| = sqrt( 3^2 + 5^2 ) |v1| = sqrt( 9 + 25 ) |v1| = sqrt( 34 )

jimthompson5910 (jim_thompson5910):

I'll let you do |v2|

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

so then the |v2| will equal sqrt65

jimthompson5910 (jim_thompson5910):

yep, you got it

OpenStudy (anonymous):

How would we then lable the v1 and v2 coordinates and the unit vectors on a graph? I have the unit vectors but not forsure how to put on a graph

jimthompson5910 (jim_thompson5910):

to graph v1, you just plot the point (3, 5) and draw an arrow from (0,0) to (3,5) the tip of the arrow is pointing to (3,5)

jimthompson5910 (jim_thompson5910):

do the same for v2

OpenStudy (anonymous):

Ok but how would we graph the unit vectors of (3/sqrt34) and (5/sqrt34)

jimthompson5910 (jim_thompson5910):

well you can convert that to decimal form to get a better idea of what it looks like

OpenStudy (anonymous):

Ok, but the x would be the first part and the y would be the 5///

jimthompson5910 (jim_thompson5910):

yes, x = 3/sqrt(34) = 3/5.83095 = 0.5144959

OpenStudy (anonymous):

I got 0.51 and 0.86

jimthompson5910 (jim_thompson5910):

y = 5/sqrt(34) = 5/5.83095 = 0.8574932

jimthompson5910 (jim_thompson5910):

close enough, the unit vector pointing in the same direction as v1 is < 0.51, 0.86 >

OpenStudy (anonymous):

The 0.51=x the 0.86=y?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Alright thank yoU!

OpenStudy (anonymous):

If i could give youanother metal i would!!

jimthompson5910 (jim_thompson5910):

you're welcome and it's fine